What is Doppler Effect?

       A police radar set emits a parallel beam of electromagnetic radiation at wavelength λo and velocity c, Which falls on a motor car moving directly towards the set with a velocity u.

       (a.) Derive an expression for the time it takes for the wave front of the radiation initially a distance

       λo from the car, to reach the car.


1
Expert's answer
2021-03-14T19:16:29-0400

The Doppler effect (or the Doppler shift) is the change in frequency of a wave in relation to an observer who is moving relative to the wave source.

"f_r=f_t(\frac{1+v/c}{1-v/c}) 
\\ f_r=f_t(\frac{c+v}{c-v})
\\f_d=f_r-f_t=2v\frac{f_t}{c-v}
\\f_d\approx 2v\frac{f_t}{c}"

"v=\frac{\lambda_0}{t}
\\f_d\approx 2v\frac{f_t}{c}\approx 2\frac{\lambda_0f_t}{tc}
\\t\approx 2\frac{\lambda_0f_t}{cf_d}"